Consider a data set containing the following values: 67, 94, 82, 80, 97, 99, 78, 91. The mean of the given value is 86. The squares of the scores are calculated as follows: 4,489, 8,836, 6,724, 6,400, 9,409, 9,801, 6,084, 8,281. If the preceding scores are sample data, the sum of squares are? If the preceding scores are population data, the SS is? If preceding scores in a sample data, the simple standard deviation is? and the sample variance is? If the preceding scors are population data, the population standard deviation is? and the population variance is?
Answer #1
S.No X1 X12
1 67.00 4489.0000
2 94.00 8836.0000
3 82.00 6724.0000
4 80.00 6400.0000
5 97.00 9409.0000
6 99.00 9801.0000
7 78.00 6084.0000
8 91.00 8281.0000
ΣX 688 ΣX2 60024.0000

1)If the preceding scores are sample data, the sum of squares are =\sumx2-(\sumx)2/n=60024-6882/8

=856

2) If the preceding scores are population data, the SS is\sumx2-(\sumx)2/n= 856

3)

If preceding scores in a sample data, the simple standard deviation is =(SS/(n-1))1/2 =11.058

4) sample variance =SS/(n-1) =122.2857

5) If the preceding scors are population data, the population standard deviation=(SS/(n)1/2 =10.344

6) population variance is =SS/n =107